# Semantic Component Analysis: Discovering Patterns in Short Texts Beyond Topics

### v2

@@ 6 Conclusion
We introduce Semantic Component Analysis as a novel topic modeling method that can discover multiple semantic components per document. We show that the components found by SCA indeed provide new, more nuanced insights into the data while maintaining a noise rate close to zero and matching the state-of-the-art in terms of topic quality. Furthermore, we show that the method is able to generalize to multiple languages, including a low-resource African language, and to provide a faithful transformation of the embeddings into the component space. We believe that SCA provides a substantial improvement in large-scale text analysis and opens up new possibilities for the interpretation of semantic patterns in short texts.
@@
While SCA provides a novel perspective on topic modeling by providing an efficient way to discover multiple, nuanced semantic components per sample and thus a more detailed explanation of the pattern distribution in a dataset, it has its limitations. The method is not able to capture non-linear, multi-dimensional components in the data, as it is based on a linear decomposition of the embedding space. In other words, SCA is based in the linear representation hypothesis Park et al. (2024). To some extent, the iterative procedure of SCA is able to mitigate this problem: Because each cluster is represented by the centroid, the decomposition might retain a part of the cluster-information in the residual embeddings in cases, when one linear unit vector is not able to capture the full complexity in terms of dimension or linearity. In such cases, similar components might be found in subsequent iterations, which can be merged by the overlap threshold.
Furthermore, while we show SCA’s applicability to a range of datasets and languages, we restrict our investigation to short texts, more precisely Tweets. We argue that problems with longer documents can partially be mitigated by splitting them into smaller parts, such as sentence level, and apply SCA to the resulting dataset. Since SCA scales to large numbers of samples even on small systems, this should be feasible for most applications.
As is common with topic modeling, the method’s token-based representation of topics has its limitations, too, as it might not able to capture the full complexity of the semantic components found. While additional methods like LLM summarizations and Medoid representations can provide additional insights, a core problem with more expressive approaches are uncertain faithfulness, a lack of good evaluation metrics as well as increased computational cost.
@@
Firstly, there is a risk of misinterpretation or misrepresentation of the discovered semantic components. The interpretation of these components relies on the assumption that they represent meaningful and coherent topics or themes in the data. However, there is always a possibility of misinterpreting or misrepresenting the true meaning of these components, leading to incorrect conclusions or biased interpretations. Even earlier in the pipeline, there is a risk of algorithmic bias in the clustering and representation process. Especially the base embedding model can introduce biases based on its training data and model architecture that SCA can inherit. It is important to carefully select and validate the base embedding model to ensure as fair and unbiased representations as possible.
Secondly, there is a risk of privacy infringement when working with sensitive or personal data. SCA relies on analyzing and clustering text data, which may contain sensitive information about individuals or groups. It is important to handle and protect this data in accordance with privacy regulations and ethical guidelines and to ensure that the semantic components extracted from the data do not reveal sensitive or personally identifiable information to unauthorized parties. The usage of the data in our evaluation does not expose any individual to such risks, as the data was either publicly available to begin with (the Hausa Dataset and Trump Archive) or does not contain any data produced by private Twitter accounts, as the Chinese News Data we scrape only consists of Tweets by the official accounts of the news outlets we include.
Thirdly, the nature of online social media data, such as Tweets, can introduce additional risks and challenges. Social media data can be subject to manipulation or misinformation, and contain offensive or harmful content. SCA may inadvertently amplify or propagate such content if not properly filtered or evaluated. Additionally, using the insights from SCA might inadvertently lead to amplifying or reinforcing existing biases or stereotypes. It is important to consider the potential impact of the data and the extracted semantic components on individuals, communities, and society as a whole.
@@ A.1 Hyperparameter Discussion
Our method involves two hyperparameters that influence the decomposition step, $\alpha$ and $\mu$. Intuitively, both can be used to influence the level of superposition allowed in the decomposition, i.e. the number of semantic components that can occupy linearly dependent parts of the embedding space and thus the the level of detail and the number of discovered semantic components. This can be necessary in cases of large data where the number of semantic components assumed is higher than the dimension of the embedding space.
@@
Furthermore, we introduce the merging threshold $\theta$ to merge components that are similar in terms of their token representation. This is necessary to avoid duplicate components that arise for large clusters spread across a non-linear or multi-dimensional part of the embedding space. We note, that the choice of $\theta$ is somewhat arbitrary and should be chosen based on the specific use case. In our evaluation, we set $\theta=0.6$ for the Trump dataset, which resulted in 11 components being merged. For the Chinese and Hausa datasets, we did not employ component merging.
@@ A.2 SCA as a Transformation
The decomposition step in SCA is a linear transformation of the embeddings into a space where each dimension corresponds to the score along one semantic component. The decomposition is based on the cosine similarity between the embedding and the centroid of the cluster, which is the representation of the semantic component. For an embedding $\mathbf{x}$ of a sample $s$, e define the $i$-th residual in an inductive way:
@@
where $\mathbf{v}_{i}$ is the $i$-th component and $\alpha_{i,i}=\frac{\langle\mathbf{x}_{i},\mathbf{v}_{i}\rangle}{||\mathbf{x}_{i}||}$. The activation of the sample $\mathbf{x}$ along the $j$-th is thus defined as the dot product of the residual and the component:
@@
From this, we can assign additional components to the samples by computing the activation along the components and applying a threshold per iteration to choose the relevant ones. This ensures, that the decomposition is able to capture multiple semantic components per sample even beyond the cluster assignments.
@@ A.3 Stopping Criteria
  • F: Terminate after a fixed number of iterations $I$
@@ B.1 Implementation and Preprocessing
B.2 Evaluation Metrics
@@
  • Noise rate: The fra¥ction of documents that are not assigned to any cluster in the first iteration (1st) and over all iterations.
@@
  • Top 10 topic diversity: As defined in Terragni et al. (2021), it is the number of unique Top-10 representative tokens (as computed by c-TF-IDF weighting) divided by the total number of representative tokens.
@@
from this, we define additional metrics to quantify the novelty of the semantic components uncovered by SCA compared to the topics found by BERTopic:
@@ C.1 Run Statistics and Grid Run
C.2 Trump Dataset: Components
@@
Figure 5: Iteration 1: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids.
@@
Figure 6: Iteration 2: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. The overlap scores refer to the maximum token overlap (top) with any component from the first iteration, i.e. a BERTopic run, and the maximum sample overlap (bottom) with any cluster in the cluster hierarchy of HDBSCAN from the first iteration (i.e. all possible topics that could be found through BERTopic).
@@
Figure 7: Iteration 4: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. The overlap scores refer to the maximum token overlap (top) with any component from the first iteration, i.e. a BERTopic run, and the maximum sample overlap (bottom) with any cluster in the cluster hierarchy of HDBSCAN from the first iteration (i.e. all possible topics that could be found through BERTopic).
C.3 Chinese News Outlet Tweets: Components
Figure 8: Iteration 1: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids.
Figure 9: Iteration 2: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
Figure 10: Iteration 3: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
C.4 Hausa Dataset: Components
@@
Figure 11: Iteration 1: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids.
@@
Figure 12: Iteration 2: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
@@
Figure 13: Iteration 3: Top-10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
C.5 Performance of Transformed Embeddings on MTEB

### v3

@@ 6 Conclusion
We introduce Semantic Component Analysis as a novel topic modeling method that can discover multiple topics and scale to large datasets. We show that the topics found by SCA indeed provide new, more nuanced insights into the data while maintaining a noise rate close to zero and matching the large-scale SOTA in terms of topic quality. Furthermore, we show that the method is able to generalize to multiple languages. However, future work is required to address the problem of generic, overpredicted clusters and the interpretability of the topic representations, perhaps by downweighing larger clusters and exploring other forms of cluster representations respectively. Taken together, we believe that SCA provides a substantial improvement in large-scale short text analysis.
@@
While SCA provides a novel perspective on topic modeling by providing an efficient way to discover multiple topics per sample and thus a more detailed explanation of the pattern distribution in a dataset, it has its limitations.
As is common with pipelines that rely on clustering embeddings, SCA inherits several potential limitations inherent to this class of algorithms. Firstly, using a sequence of modular steps might introduce errors that could be propagated and thus amplified in subsequent steps. For example, if the clustering step fails to find meaningful clusters, the topics found will likely be of low quality. To an extent, this problem can be mitigated by carefully chosing the modules and hyperparameters in accordance with the data at hand and performing multiple runs to test robustness. However, the risk of error propagation in clustering-based topic modelling remains and should be investigated further in future work.
Furthermore, the similarity modeled by the input embeddings is crucial for the quality of the topics found. If the embedding model does not capture the semantic similarity in a way that is suitable for the data at hand, the topics found will likely be of low quality. This is especially relevant in low-resource languages or domains, where high-quality embedding models are often not available. While SCA can be used with any embedding model, it is important to carefully select and validate the base embedding model to ensure as fair and unbiased representations as possible.
Besides, SCA is not able to capture non-linear, multi-dimensional components in the data, as it is based on a linear decomposition of the embedding space. In other words, SCA is based in the linear representation hypothesis Park et al. (2024). To some extent, the iterative procedure of SCA is able to mitigate this problem: Because each cluster is represented by the centroid, the decomposition might retain a part of the cluster-information in the residual embeddings in cases, when one linear unit vector is not able to capture the full complexity in terms of dimension or linearity. In such cases, similar topics might be found in subsequent iterations, which can be merged by the overlap threshold $\theta$ to avoid duplicate topics.
Furthermore, while we show SCA’s applicability to a range of datasets and languages, we restrict our investigation to short texts, more precisely Tweets. We argue that problems with longer documents can partially be mitigated by splitting them into smaller parts, such as paragraph level, and apply SCA to the resulting dataset. Since SCA scales to large numbers of samples even on small systems, this should be feasible for most applications. Adding to this, our evaluation does not include a user study to assess the interpretability of the topics found. While we provide qualitative examples and find that the topics discovered match the expected distribution in the datasets, a user study could provide additional insights into the practical applicability of SCA.
As is common with topic modeling, the method’s token-based representation of topics has its limitations, too, as it might not able to capture the full complexity of the topics found. While additional methods like LLM summarizations and Medoid representations can provide additional insights, a core problem with more expressive approaches are uncertain faithfulness, a lack of good evaluation metrics as well as increased computational cost.
@@
Firstly, there is a risk of misinterpretation or misrepresentation of the discovered topics. The interpretation of these topics relies on the assumption that they represent meaningful and coherent topics or themes in the data. However, there is always a possibility of misinterpreting or misrepresenting the true meaning of these topics, leading to incorrect conclusions or biased interpretations. Even earlier in the pipeline, there is a risk of algorithmic bias in the clustering and representation process. Especially the base embedding model can introduce biases based on its training data and model architecture that SCA can inherit. It is important to carefully select and validate the base embedding model to ensure as fair and unbiased representations as possible.
Secondly, there is a risk of privacy infringement when working with sensitive or personal data. SCA relies on analyzing and clustering text data, which may contain sensitive information about individuals or groups. It is important to handle and protect this data in accordance with privacy regulations and ethical guidelines and to ensure that the topics extracted from the data do not reveal sensitive or personally identifiable information to unauthorized parties. The usage of the data in our evaluation does not expose any individual to such risks, as the data was either publicly available to begin with (the Hausa Dataset and Trump Archive) or does not contain any data produced by private Twitter accounts, as the Chinese News Data we scrape only consists of Tweets by the official accounts of the news outlets we include.
Thirdly, the nature of online social media data, such as Tweets, can introduce additional risks and challenges. Social media data can be subject to manipulation or misinformation, and contain offensive or harmful content. SCA may inadvertently amplify or propagate such content if not properly filtered or evaluated. Additionally, using the insights from SCA might inadvertently lead to amplifying or reinforcing existing biases or stereotypes. It is important to consider the potential impact of the data and the extracted topics on individuals, communities, and society as a whole.
@@
Use of AI assistants.
The authors acknowledge the use of ChatGPT for correcting grammatical errors and improving the overall readability of the paper. Furthermore, GitHub Copilot was used to assist in writing the code. The authors have reviewed and edited the output of these AI assistants to ensure accuracy and clarity.
@@ A.1 Hyperparameter Discussion
Our method involves two hyperparameters that influence the decomposition step, $\alpha$ and $\mu$. Intuitively, both can be used to influence the level of superposition allowed in the decomposition, i.e. the number of semantic components that can occupy linearly dependent parts of the embedding space and thus the the level of detail and the number of discovered topics. This can be necessary in cases of large data where the number of topics assumed is higher than the dimension of the embedding space.
@@
Furthermore, we introduce the merging threshold $\theta$ to merge components that are similar in terms of their token representation. This is necessary to avoid duplicate components that arise for large clusters spread across a non-linear or multi-dimensional part of the embedding space. We note that the choice of $\theta$ is somewhat arbitrary and should be chosen based on the specific use case. In our evaluation, we set $\theta=0.6$ for the Trump dataset, which resulted in 11 components being merged. For the Chinese and Hausa datasets, we did not employ component merging.
@@ A.2 SCA as a Transformation
The decomposition step in SCA is a linear transformation of the embeddings into a space where each dimension corresponds to the score along one semantic component. The decomposition is based on the cosine similarity between the embedding and the centroid of the cluster, which is the semantic component representing the connected topic. For an embedding $\mathbf{x}$ of a sample $s$, e define the $i$-th residual in an inductive way:
@@
[…]
@@
From this, we can assign additional components to the samples by computing the activation along the components and applying a threshold per iteration to choose the relevant ones. This ensures, that the decomposition is able to capture multiple topics per sample even beyond the cluster assignments.
@@ A.3 Stopping Criteria
  • F: Terminate after a fixed number of iterations II
@@ B.1 Implementation and Preprocessing
B.2 TopicGPT Comparison
We compare SCA against the TopicGPT method Pham et al. (2024), which uses a large language model to generate topics from the embeddings. To ensure a fair comparison in the resource-restricted scenario SCA is targeted, we aim to evaluate TopicGPT with backbone models of similar size as the embedding model we use for our experiments, i.e. 278M paramaters for paraphrase-multilingual-mpnet-base-v2. We use the instruction tuned Llama-3.2 model Grattafiori (2024) with 1B and 8B parameters, namely Llama-3.2-1B-Instruct and Llama-3.2-8B-Instruct and run TopicGPT in the standard configuration without changing any hyperparameters or steps of the procedure as defined in their GitHub repository. [^5] See https://github.com/chtmp223/topicGPT
We note that the authors of TopicGPT report their results using the GPT-3.5 model through the OpenAI API, which is not available for local installations. We include these results in Table 1 for comparison.
Furthermore, note the extensive compute budget necessary to run TopicGPT on the Bills dataset, which has only 32,661 samples. There, the authors report a total cost of \$88 US dollars, of which \$30 is for the topic generation, \$10 for topic refinement and \$48 for the for the assignment and correction step using the OpenAI API. Assuming the former two do not scale with dataset size, the cost for the assignment and correction step would be \$1.47 per 1000 samples. Extrapolating this cost to the Trump dataset, which has 60,053 samples, would result in a total cost of \$88.57 for the assignment and correction step alone. For the Hausa dataset, which has 511,523 samples, the cost would be \$751.75 and for the Chinese News dataset, which has 1,463,150 samples, the cost would be \$2150.30. We argue that this kind of cost is not feasible for most applications, especially in the low-resource setting SCA is targeted at.
B.3 Evaluation Metrics
@@
  • Noise rate: The fraction of documents that are not assigned to any cluster in the first iteration (1st) and over all iterations.
@@
  • TOP10 topic diversity: As defined in Terragni et al. (2021), it is the number of unique TOP10 representative tokens (as computed by c-TF-IDF weighting) divided by the total number of representative tokens.
@@
from this, we define additional metrics to quantify the novelty of the topics uncovered by SCA compared to those found by BERTopic:
@@ C.1 Run Statistics and Grid Run
C.2 Trump Dataset: SCA Topics
@@
Figure 5: Iteration 1: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids.
@@
Figure 6: Iteration 2: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids. The overlap scores refer to the maximum token overlap (top) with any component from the first iteration, i.e. a BERTopic run, and the maximum sample overlap (bottom) with any cluster in the cluster hierarchy of HDBSCAN from the first iteration (i.e. all possible topics that could be found through BERTopic).
@@
Figure 7: Iteration 4: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids. The overlap scores refer to the maximum token overlap (top) with any component from the first iteration, i.e. a BERTopic run, and the maximum sample overlap (bottom) with any cluster in the cluster hierarchy of HDBSCAN from the first iteration (i.e. all possible topics that could be found through BERTopic).
C.3 Trump Dataset: TopicGPT Topics
The TopicGPT baseline does not perform well with a backbone model (Llama-3.2-1B) of comparable size to our embedding model (278M parameters). In particular, the topic generation phase seems to have troubles identifying meaningfull patterns. We plot the resulting topics in Figure 8.
Topic ID Topic Description
Trade The exchange of goods, services, and capital.
Agriculture The production, cultivation, and harvesting of crops and livestock.
Fluopyram A medication used to treat certain types of pain.
Ballot Scam A type of election-related scam where voters are tricked into returning
a ballot to the wrong address.
None General Topics
Figure 8: Topics generated by TopicGPT on the Trump dataset.
C.4 Chinese News Outlet Tweets: Components
Figure 9: Iteration 1: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids.
Figure 10: Iteration 2: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
Figure 11: Iteration 3: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
C.5 Hausa Dataset: Components
@@
Figure 12: Iteration 1: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids.
@@
Figure 13: Iteration 2: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
@@
Figure 14: Iteration 3: TOP10 components by number of samples N together with their c-TF-IDF representations and medoids. Overlap describes the maximum number of tokens a component’s representation has in common with any component from the first iteration.
C.6 Performance of Transformed Embeddings on MTEB
@@
C.7 Example from Pham et al. (2024)
Pham et al. (2024) provide an example topic assignment for a sample from the Bills dataset to illustrate the qulitative validity of their approach. Looking at the topic assigned by SCA, we argue that our method is also able to match the correct topic relating to education, see 3.
Ground truth Sample SCA LDA TopicGPT
Education Bills Perkins Fund for Equity and Excellence. This bill amends the Carl D. Perkins Career and Technical Education Act of 2006 to replace the existing Tech Prep program with a new competitive grant program to support career and technical education. Under the program, local educational agencies and their partners may apply for grant funding to support: career and technical education programs that are aligned with postsecondary education programs, dual or concurrent enrollment programs and early college programs, certain evidence-based strategies and delivery models related to career and technical education, teacher and leader experiential […] leas, students, student, school, schools, ihes, 1965, education, elementary, academic City infrastructure: city, building, area, new, park Education: Mentions policies and programs related to higher education and student loans.
Table 3: Example from the Bills dataset from Pham et al. (2024). The SCA method is able to assign the correct topic relating to education. Results marked with are copied from Pham et al. (2024).